OpenStudy (anonymous):

Part A: The sun produces 3.9 ⋅ 1033 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 2.45 ⋅ 105 seconds? Part B: Which is the more reasonable measurement of the distance between the tracks on a DVD: 7.4 ⋅ 10−4 mm or 7.4 ⋅ 104 mm? Justify your answer.

OpenStudy (anonymous):

\[3.9\times 10^{33}\times 2.25\times 10^5\]

OpenStudy (anonymous):

add the exponents

OpenStudy (anonymous):

and multiply \(3.9\times 2.45\)

OpenStudy (anonymous):

9.56 x 10^38?